# Parameter Efficient Continual Learning with Dynamic Low-Rank Adaptation

### v3

@@ 4 Results
Table 2: Class-IL performance comparison on ImageNet-R across 5, 10, and 20 task sequences. Baseline results with standard deviation are taken from [liang2024inflora] and that of PEARL are averaged over 3 runs with errors indicating standard deviation across these runs. The rest of the results are taken from their respective manuscripts. Best results are highlighted in bold. SeqLoRA is a sequential learning baseline which fine-tunes LoRA components on multiple tasks without any additional support.
@@
L2P [wang2022learning] CVPR 2022 64.13 $\pm$ 0.78 68.66 $\pm$ 0.41 62.54 $\pm$ 0.24 67.98 $\pm$ 0.27 57.92 $\pm$ 0.28 64.57 $\pm$ 0.29
DualPrompt [DBLP:conf/eccv/0002ZESZLRSPDP22] ECCV 2022 67.88 $\pm$ 0.17 71.16 $\pm$ 0.31 65.41 $\pm$ 0.52 69.39 $\pm$ 0.43 61.00 $\pm$ 0.72 65.80 $\pm$ 0.67
CODA-P [smith2023coda] CVPR 2023 73.09 $\pm$ 0.21 76.91 $\pm$ 0.21 71.47 $\pm$ 0.35 75.82 $\pm$ 0.29 67.28 $\pm$ 0.30 72.34 $\pm$ 0.17
C-LoRA [DBLP:journals/corr/abs-2304-06027] CoRR 2023 75.85 $\pm$ 0.31 78.85 $\pm$ 0.34 71.89 $\pm$ 0.45 75.33 $\pm$ 0.28 65.71 $\pm$ 0.60 70.63 $\pm$ 0.85
LAE [gao2023unified] ICCV 2023 73.84 $\pm$ 0.14 77.29 $\pm$ 0.45 71.70 $\pm$ 0.39 76.71 $\pm$ 0.10 66.98 $\pm$ 0.35 73.72 $\pm$ 0.05
InfLoRA [liang2024inflora] CVPR 2024 77.52 $\pm$ 0.37 82.01 $\pm$ 0.12 75.65 $\pm$ 0.14 80.82 $\pm$ 0.24 71.01 $\pm$ 0.45 77.28 $\pm$ 0.45
EASE [zhou2024expandable] CVPR 2024 70.80     - 73.37     - 78.80     -
SEMA [SEMA] CVPR 2025 79.78 84.75 78.00 83.56 74.53 81.75
DIA [DIA] CVPR 2025     -     - 79.03 85.61 76.32 83.51
TSVD [TSVD] ICLR 2025 73.58     - 74.55     - 79.13     -
@@
Comparison with PEFT CL approaches: As PEARL entails dynamic allocation of LoRA components, we compare PEARL with state-of-the-art PEFT CL methods across 5, 10, and 20 tasks in ImageNet-R [hendrycks2021many], in Table 2. The experiment is conducted using the ViT-B/16 backbone [dosovitskiy2020image] pre-trained on ImageNet-21k. As can be seen, PEARL (ViT) shows a significant improvement in both incremental and final accuracy, with absolute gains ranging from 9% to 14%. The performance improvement is more pronounced in 5 task sequence and slightly decreases thereafter as we move towards 10 and then to 20 task sequences. Among considered baselines, InfoLoRA is the closest work in terms of design and performance, but falls short of PEARL (ViT)’s performance by a huge margin across all scenarios. InfLoRA injects a small number of parameters to re-parameterize the pre-trained weights and shows that finetuning these injected parameters is equivalent to finetuning the pre-trained weights within a subspace. Since the reduction matrix $B$ infLoRA is handcrafted rather than being learnt from the training data, InfLoRA suffers a performance degradation compared to PEARL (ViT). Additional comparisons between our method and adapter-based as well as prompt-based approaches can be found in Table 14 and Table 14.
@@
With and without re-initialization: The PEARL framework entails creation of task-specific, layer-specific LoRA components whose weights are initialized from SV-decomposition of task vectors. As task vectors represent the additional information on top of information contained in the reference task, one would expect LoRA components to benefit from such an initialization. Table 4 compares the effects of weight re-initialization of LoRA components, BN layers (if any), and classification layers following SV-decomposition in PEARL. In the Without re-init scenario, the weights of the LoRA components are initialized directly from the SV-decomposition of task vectors and are then fine-tuned for the current task. Conversely, in PEARL, the weights of the LoRA components, as well as the batch normalization (BN) and classification layers, are re-initialized after the creation of the LoRA components.The results indicate that weight re-initialization leads to marginally better performance for ResNet-18 on Seq-CIFAR100 (5T), statistically significant performance improvements for ViT on ImageNet-R (10T), and a remarkable two-fold increase in performance for BSC on Seq-TinyImageNet (10T). Overall, the evidence suggests that weight re-initialization positively influences PEARL across various CL architectures and settings. We speculate that fresh start due to weight re-initialization helps in avoiding local minima and thus better generalization across tasks.
Dataset Variant Metric Without re-init With re-init (Ours)
Seq-CIFAR100 (5T) R18 $A_{T}$ (%) 51.34 $\pm$ 0.71 51.97 $\pm$ 0.93
$\bar{A}$ (%) 62.37 $\pm$ 1.20 63.51 $\pm$ 0.75
Seq-TinyImageNet (10T) BSC $A_{T}$ (%) 18.71 $\pm$ 0.79 36.45 $\pm$ 0.54
$\bar{A}$ (%) 29.25 $\pm$ 0.77 48.64 $\pm$ 0.43
ImageNet-R (10T) ViT $A_{T}$ (%) 78.71 $\pm$ 0.38 83.51 $\pm$ 0.47
$\bar{A}$ (%) 84.79 $\pm$ 0.63 88.30 $\pm$ 0.48
Dynamic rank allocation across layers: PEARL leverages reference task weights and adaptively determines the rank of task-specific, layer-specific LoRA components based on the proximity of the current task to the reference task weights in the parameter space. Figure 1 illustrates the dynamic rank allocation across layers for convolutional variants for the final task in Seq-CIFAR100 (5T) and that of ViT for the final task in Imagenet-R (10T) can be found in Figure 4. From left to right, Figure 1 presents a detailed overview of the layer-wise dynamic thresholds established based on task proximity, along with the corresponding selected rank and actual rank for our convolutional variants. Consistent with the actual rank, our framework allocates lower ranks to earlier layers and higher ranks to later layers. The dynamic rank allocation is also consistent with established conventions in CNN literature, which suggest that earlier layers capture more generic features, while later layers focus on task-specific information. As a result, earlier layers require less task-specific adaptation, leading to lower rank allocations. Conversely, as we approach the later layers, more resources are necessary to adapt the reference task information to the current task, thereby justifying the higher rank allocation in these layers.
5 Conclusion
Appendix A Additional Information
A.1 Overview of experiments
Table 5: Summary of experiments.
- BSC - Seq-TinyImageNet (5T, 10T) - Task-IL - NECIL - $\bar{A}$ (%)
- Param isolation based - $A_{T}$ (%)
A.2 Justification for considered baselines
To facilitate extensive evaluation of PEARL, we consider a comprehensive array of baseline methods: (i) exemplar-based approaches that involve information reuse and task-specific adapters (e.g., TAMiL, AGILE); (ii) NECIL approaches (e.g., PASS, FeTriL, FeCAM); (iii) parameter-isolation approaches that operate within a fixed capacity (e.g., NISPA) and those that expand beyond it (e.g., PNN, PAE); and (iv) PEFT approaches for continual learning (e.g., L2P, C-LoRA, InfLoRA). Firstly, inspired by the GWT, exemplar-based baselines entail a shared network for information reuse and task-specific attention modules that re-project the representation from the common representation space to the global workspace. Likewise, the PEARL framework includes a shared reference task network for information reuse and task-specific, layer-specific low-rank adapters for orienting towards the current task. Secondly, PEARL is a rehearsal-free framework; it does not store previous task samples either directly or indirectly through generative networks. Therefore, it aligns well with the NECIL benchmark and associated approaches. Thirdly, we note that low-rank adaptation introduces task-specific parameters that are trained on the corresponding task and kept frozen afterwards, akin to parameter-isolation approaches. As PEARL exhibits partial parameter isolation with a shared reference task network and low-rank adapters, we also provide a comparison with parameter isolation approaches. Finally, since PEARL readily utilizes LoRA [hu2021lora], a prominent PEFT technique, we offer an in-depth comparison with PEFT continual learning approaches. In addition, we also provide a discussion on related dynamic low-rank adaptation techniques beyond continual learning in Section A.3.
A.3 Discussion on PEFT LoRA counterparts
PEFT CL approaches with static rank allocation: Among PEFT continual learning approaches, InfLoRA [liang2024inflora] and C-LoRA [DBLP:journals/corr/abs-2304-06027] are prominent techniques that involve low-rank adaptation of pre-trained networks, similar to PEARL. Each of these methods are tailored for CL scenarios and addresses unique challenges. InfLoRA aims to minimize the interference of new tasks on previously learned ones, thereby improving knowledge retention. However, its reliance on a handcrafted reduction matrix $B$, rather than a data-driven learned representation, limits its flexibility. C-LoRA, on the other hand, depends on a pre-trained model as its foundation, making it reliant on the availability of such models for its application. It applies LoRA to both the key and query attention mechanisms, ensuring efficient adaptation for new tasks. Additionally, C-LoRA incorporates a self-regularization strategy, which penalizes updates to parameters already modified by previous tasks. This mechanism helps preserve prior knowledge and enhances stability across sequential tasks. Despite its relative success, caution is warranted when training C-LoRA on large task sequences, as this approach suffers from scalability issues. In line with standard practices in continual learning benchmarks, we provide results for SeqLoRA, a sequential finetuning baseline without any safeguards against catastrophic forgetting. SeqLoRA experiences significant forgetting, underscoring the difficulty of maintaining performance across tasks in continual learning.
PEFT approaches with dynamic rank allocation: Parameter efficient fine-tuning techniques such as LoRA distribute trainable parameters equally across all layers, which may reduce efficiency when layers vary in their contribution towards task adaptation. To address this, adaptive low-rank adaptation strategies have emerged that dynamically adjust parameter budgets to improve fine-tuning efficiency, similar to PEARL. AdaLoRA [zhang2023adaptive] adaptively allocates the parameter budget among weight matrices according to their importance score. The importance score is computed by parameterizing the incremental updates in the form of SV-decomposition and the singular values of unimportant updates are pruned. DyLoRA [valipour-etal-2023-dylora] proposed a dynamic low-rank adaptation in which LoRA blocks are trained for a range of ranks instead of a single rank by sorting the representation learned by the adapter module at different ranks during training. More recent methods, such as DoRA [mao-etal-2024-dora] and Dynamic Rank-Selective LoRA [lu2024adaptive], have emerged, with DoRA decomposing high-rank LoRA layers into structured single-rank components, while Dynamic Rank-Selective LoRA adaptively assigns ranks to LoRA modules based on their relevance to the current data in CL in vision language models. Although these approaches address a similar challenge as PEARL, these are not tailored for CL in custom, compact models.
A.4 Limitations and Future work
We proposed PEARL, a versatile parameter efficient CL framework with dynamic low-rank adaptation. PEARL is generic and can be effectively applied to both convolutional neural networks and Vision Transformers. We provide extensive analysis to exhibit the effectiveness of PEARL across multitude of datasets, settings and scenarios in CL. However, PEARL has several limitations: Firstly, as a parameter isolation approach, PEARL scales with the increasing number of tasks. Although this growth is limited, it still results in additional parameters for each new task, leading to a larger model. Moreover, with every new task, each image has to pass through every sub-network for Class-IL inference, resulting slower inference. Secondly, PEARL assumes knowledge of task boundaries, which are not always available in real-world scenarios. This assumption is common among parameter isolation approaches, as new parameters are typically introduced at the onset of each task. Thirdly, PEARL requires fine-tuning a full pre-trained model on the current task for several epochs. As model sizes grow, this may become impractical on resource-constrained devices. In our future work, we aim to explore alternatives for defining task boundaries and efficient estimation of task vectors without the need to fine-tune the entire pre-trained model.
Appendix B Additional Analysis
@@
B.1 Comparison with fixed rank as a baseline
The PEFT CL approaches offer lightweight training schedules while effectively retaining the consolidated knowledge of pre-trained models. However, LoRA components in these methods are particularly sensitive to rank selection, which can lead to suboptimal resource allocation and diminished performance [valipour2022dylora]. Table 6 summarizes the average model size and performance metrics across various ranks for LoRA components. As the rank increases, performance improves at the expense of increased model size; however, beyond a certain rank, the performance gains do not really justify the additional increase in model size. Consequently, tuning the rank is essential to strike the right balance in this trade-off where model growth remains modest with optimum performance. To this end, PEARL employs a task-specific, layer-specific dynamic rank allocation enabling an optimal trade-off between model size and performance across different datasets and architectures.
B.2 Comparison with convolutional parameter isolation approaches
@@
Figure 2: Task-IL performance comparison with parameter isolation approaches on Seq-CIFAR100 (20T). We report the final Task-IL accuracies of each task after training on all tasks. Methods compared are NISPA [gurbuz2022nispa], PNN [rusu2016progressive], PAE [hung2019increasingly], CPG [hung2019compacting], PackNet [mallya2018packnet], and CLNP [golkar2019continual]
B.3 Comparison with complete parameter isolation

### v4

@@ 4.1 Comparison with convolutional CL approaches
Table 2: Class-IL performance comparison on ImageNet-R across 5, 10, and 20 task sequences. Baseline results with standard deviation are taken from Liang & Li (2024) and that of PEARL are averaged over 3 runs with errors indicating standard deviation across these runs. The rest of the results are taken from their respective manuscripts. Best results are highlighted in bold. SeqLoRA is a sequential learning baseline which fine-tunes LoRA components on multiple tasks without any additional support.
@@
L2P (Wang et al., 2022e) CVPR 2022 64.13 $\pm$ 0.78 68.66 $\pm$ 0.41 62.54 $\pm$ 0.24 67.98 $\pm$ 0.27 57.92 $\pm$ 0.28 64.57 $\pm$ 0.29
DualPrompt (Wang et al., 2022d) ECCV 2022 67.88 $\pm$ 0.17 71.16 $\pm$ 0.31 65.41 $\pm$ 0.52 69.39 $\pm$ 0.43 61.00 $\pm$ 0.72 65.80 $\pm$ 0.67
CODA-P (Smith et al., 2023b) CVPR 2023 73.09 $\pm$ 0.21 76.91 $\pm$ 0.21 71.47 $\pm$ 0.35 75.82 $\pm$ 0.29 67.28 $\pm$ 0.30 72.34 $\pm$ 0.17
C-LoRA (Smith et al., 2023a) CoRR 2023 75.85 $\pm$ 0.31 78.85 $\pm$ 0.34 71.89 $\pm$ 0.45 75.33 $\pm$ 0.28 65.71 $\pm$ 0.60 70.63 $\pm$ 0.85
LAE (Gao et al., 2023b) ICCV 2023 73.84 $\pm$ 0.14 77.29 $\pm$ 0.45 71.70 $\pm$ 0.39 76.71 $\pm$ 0.10 66.98 $\pm$ 0.35 73.72 $\pm$ 0.05
InfLoRA (Liang & Li, 2024) CVPR 2024 77.52 $\pm$ 0.37 82.01 $\pm$ 0.12 75.65 $\pm$ 0.14 80.82 $\pm$ 0.24 71.01 $\pm$ 0.45 77.28 $\pm$ 0.45
EASE (Zhou et al., 2024b) CVPR 2024 70.80     - 73.37     - 78.80     -
SEMA (Wang et al., 2025) CVPR 2025 79.78 84.75 78.00 83.56 74.53 81.75
DIA (Li et al., 2025) CVPR 2025     -     - 79.03 85.61 76.32 83.51
TSVD (Peng et al., 2025) ICLR 2025 73.58     - 74.55     - 79.13     -
@@
4.2 Comparison with PEFT CL approaches
As PEARL entails dynamic allocation of LoRA components, we compare PEARL with state-of-the-art PEFT CL methods across 5, 10, and 20 tasks in ImageNet-R (Hendrycks et al., 2021), in Table 2. The experiment is conducted using the ViT-B/16 backbone (Dosovitskiy et al., 2021) pre-trained on ImageNet-21k. As can be seen, PEARL (ViT) shows a significant improvement in both incremental and final accuracy, with absolute gains ranging from 9% to 14%. The performance improvement is more pronounced in 5 task sequence and slightly decreases thereafter as we move towards 10 and then to 20 task sequences. Among considered baselines, InfoLoRA is the closest work in terms of design and performance, but falls short of PEARL (ViT)’s performance by a huge margin across all scenarios. InfLoRA injects a small number of parameters to re-parameterize the pre-trained weights and shows that finetuning these injected parameters is equivalent to finetuning the pre-trained weights within a subspace. Since the reduction matrix $B$ infLoRA is handcrafted rather than being learnt from the training data, InfLoRA suffers a performance degradation compared to PEARL (ViT). Additional comparisons between our method and adapter-based as well as prompt-based approaches can be found in Table 14 and Table 15.
@@ 4.2 Comparison with PEFT CL approaches
4.3 Comparison with fixed rank as a baseline
The PEFT CL approaches offer lightweight training schedules while effectively retaining the consolidated knowledge of pre-trained models. However, LoRA components in these methods are particularly sensitive to rank selection, which can lead to suboptimal resource allocation and diminished performance (Valipour et al., 2022). Table 4 summarizes the average model size and performance metrics across various ranks for LoRA components. As the rank increases, performance improves at the expense of increased model size; however, beyond a certain rank, the performance gains do not really justify the additional increase in model size. Consequently, tuning the rank is essential to strike the right balance in this trade-off where model growth remains modest with optimum performance. To this end, PEARL employs a task-specific, layer-specific dynamic rank allocation enabling an optimal trade-off between model size and performance across different datasets and architectures.
@@ 4.3 Comparison with fixed rank as a baseline
4.4 Comparison with convolutional parameter isolation approaches
@@
Figure 1: Task-IL performance comparison with parameter isolation approaches on Seq-CIFAR100 (20T). We report the final Task-IL accuracies of each task after training on all tasks. Methods compared are NISPA (Gurbuz & Dovrolis, 2022), PNN (Rusu et al., 2016), PAE (Hung et al., 2019b), CPG (Hung et al., 2019a), PackNet (Mallya & Lazebnik, 2018), and CLNP (Golkar et al., 2019)
4.5 With and without re-initialization
The PEARL framework entails creation of task-specific, layer-specific LoRA components whose weights are initialized from SV-decomposition of task vectors. As task vectors represent the additional information on top of information contained in the reference task, one would expect LoRA components to benefit from such an initialization. Table 5 compares the effects of weight re-initialization of LoRA components, BN layers (if any), and classification layers following SV-decomposition in PEARL. In the Without re-init scenario, the weights of the LoRA components are initialized directly from the SV-decomposition of task vectors and are then fine-tuned for the current task. Conversely, in PEARL, the weights of the LoRA components, as well as the batch normalization (BN) and classification layers, are re-initialized after the creation of the LoRA components.The results indicate that weight re-initialization leads to marginally better performance for ResNet-18 on Seq-CIFAR100 (5T), statistically significant performance improvements for ViT on ImageNet-R (10T), and a remarkable two-fold increase in performance for BSC on Seq-TinyImageNet (10T). Overall, the evidence suggests that weight re-initialization positively influences PEARL across various CL architectures and settings. We speculate that fresh start due to weight re-initialization helps in avoiding local minima and thus better generalization across tasks.
Dataset Variant Metric Without re-init With re-init (Ours)
Seq-CIFAR100 (5T) R18 $A_{T}$ (%) 51.34 $\pm$ 0.71 51.97 $\pm$ 0.93
$\bar{A}$ (%) 62.37 $\pm$ 1.20 63.51 $\pm$ 0.75
Seq-TinyImageNet (10T) BSC $A_{T}$ (%) 18.71 $\pm$ 0.79 36.45 $\pm$ 0.54
$\bar{A}$ (%) 29.25 $\pm$ 0.77 48.64 $\pm$ 0.43
ImageNet-R (10T) ViT $A_{T}$ (%) 78.71 $\pm$ 0.38 83.51 $\pm$ 0.47
$\bar{A}$ (%) 84.79 $\pm$ 0.63 88.30 $\pm$ 0.48
4.6 Dynamic rank allocation across layers
PEARL leverages reference task weights and adaptively determines the rank of task-specific, layer-specific LoRA components based on the proximity of the current task to the reference task weights in the parameter space. Figure 2 illustrates the dynamic rank allocation across layers for convolutional variants for the final task in Seq-CIFAR100 (5T) and that of ViT for the final task in Imagenet-R (10T) can be found in Figure 4. From left to right, Figure 2 presents a detailed overview of the layer-wise dynamic thresholds established based on task proximity, along with the corresponding selected rank and actual rank for our convolutional variants. Consistent with the actual rank, our framework allocates lower ranks to earlier layers and higher ranks to later layers. The dynamic rank allocation is also consistent with established conventions in CNN literature, which suggest that earlier layers capture more generic features, while later layers focus on task-specific information. As a result, earlier layers require less task-specific adaptation, leading to lower rank allocations. Conversely, as we approach the later layers, more resources are necessary to adapt the reference task information to the current task, thereby justifying the higher rank allocation in these layers.
4.7 Effect of pre-training on PEARL
Method Seq-CIFAR100 (5T) Seq-TinyImageNet (10T)
$_{T}A¯\bar{A}$% $_{T}A¯\bar{A}$%
4.8 Effect of random class ordering
The Table 7 displays the Class-IL performance on ImageNet-R (10T) under five different class orders, which differ from those presented in Table 2. In this setting, we randomize class assignments across tasks by setting different random seeds, allowing us to assess the impact of class ordering on PEARL. Although the final Class-IL accuracy ($A_{T}$%) is slightly lower than the previously reported result of $88.76\pm 0.69$, the results remain consistent and robust across these new class orders, highlighting the method’s stability. Furthermore, both incremental and final accuracies show stability across varying class orderings, reinforcing the reliability of the approach. Even across all different class orderings, PEARL (ViT) consistently outperforms the considered baselines in Table 2 by a huge margin. In conclusion, the findings indicate that PEARL maintains robust performance and generalizability across varying class orders, affirming its reliability as an effective framework for CL, even in randomized conditions.
Class Order # Params (Millions) $_{T}A¯\bar{A}$%
1 92.78 86.63 91.24
2 93.46 86.76 92.24
3 93.41 86.50 91.54
4 93.26 86.73 91.69
5 92.83 86.86 91.77
Average 93.14 86.70 91.70
4.9 Growth in model size
Dataset Backbone Method 5 Tasks 10 Tasks 20 Tasks
Seq-CIFAR100 R18 Complete Param. Isolation 55.89 111.73 223.42
PEARL 12.56 13.27 13.55
BSC Complete Param. Isolation 6.32 12.60 25.14
PEARL 1.63 1.89 2.30
ImageNet-R ViT Complete Param. Isolation 452.86 905.72 1,811.44
PEARL 92.46 93.27 95.35
5 Conclusion
Appendix A Limitations and Future work
We provided extensive analysis to exhibit the effectiveness of PEARL across multitude of datasets, settings and scenarios in CL. However, PEARL has several limitations: Firstly, as a parameter isolation approach, PEARL scales with the increasing number of tasks. Although this growth is limited, it still results in additional parameters for each new task, leading to a larger model. Moreover, with every new task, each image has to pass through every sub-network for Class-IL inference, resulting slower inference. Secondly, PEARL assumes knowledge of task boundaries, which are not always available in real-world scenarios. This assumption is common among parameter isolation approaches, as new parameters are typically introduced at the onset of each task. Thirdly, PEARL requires fine-tuning a full pre-trained model on the current task for several epochs. As model sizes grow, this may become impractical on resource-constrained devices. Finally, task vector initialization in PEARL assumes a precise measure of task proximity, but in practice, there can be multiple proximity criteria can be explored and validated. In our future work, we aim to explore alternatives for defining task boundaries and efficient estimation of task vectors without the need to fine-tune the entire pre-trained model.
Appendix B Additional Information
B.1 Overview of experiments
Table 9: Summary of experiments.
- BSC - Seq-TinyImageNet (5T, 10T) - Task-IL - NECIL - $\bar{A}$ (%)
- Param isolation based - $A_{T}$ (%)
B.2 Justification for considered baselines
To facilitate extensive evaluation of PEARL, we consider a comprehensive array of baseline methods: (i) exemplar-based approaches that involve information reuse and task-specific adapters (e.g., TAMiL, AGILE); (ii) NECIL approaches (e.g., PASS, FeTriL, FeCAM); (iii) parameter-isolation approaches that operate within a fixed capacity (e.g., NISPA) and those that expand beyond it (e.g., PNN, PAE); and (iv) PEFT approaches for continual learning (e.g., L2P, C-LoRA, InfLoRA). Firstly, inspired by the GWT, exemplar-based baselines entail a shared network for information reuse and task-specific attention modules that re-project the representation from the common representation space to the global workspace. Likewise, the PEARL framework includes a shared reference task network for information reuse and task-specific, layer-specific low-rank adapters for orienting towards the current task. Secondly, PEARL is a rehearsal-free framework; it does not store previous task samples either directly or indirectly through generative networks. Therefore, it aligns well with the NECIL benchmark and associated approaches. Thirdly, we note that low-rank adaptation introduces task-specific parameters that are trained on the corresponding task and kept frozen afterwards, akin to parameter-isolation approaches. As PEARL exhibits partial parameter isolation with a shared reference task network and low-rank adapters, we also provide a comparison with parameter isolation approaches. Finally, since PEARL readily utilizes LoRA (Hu et al., 2021), a prominent PEFT technique, we offer an in-depth comparison with PEFT continual learning approaches. In addition, we also provide a discussion on related dynamic low-rank adaptation techniques beyond continual learning in Section B.3.
B.3 Discussion on PEFT LoRA counterparts
PEFT CL approaches with static rank allocation: Among PEFT continual learning approaches, InfLoRA (Liang & Li, 2024) and C-LoRA (Smith et al., 2023a) are prominent techniques that involve low-rank adaptation of pre-trained networks, similar to PEARL. Each of these methods are tailored for CL scenarios and addresses unique challenges. InfLoRA aims to minimize the interference of new tasks on previously learned ones, thereby improving knowledge retention. However, its reliance on a handcrafted reduction matrix $B$, rather than a data-driven learned representation, limits its flexibility. C-LoRA, on the other hand, depends on a pre-trained model as its foundation, making it reliant on the availability of such models for its application. It applies LoRA to both the key and query attention mechanisms, ensuring efficient adaptation for new tasks. Additionally, C-LoRA incorporates a self-regularization strategy, which penalizes updates to parameters already modified by previous tasks. This mechanism helps preserve prior knowledge and enhances stability across sequential tasks. Despite its relative success, caution is warranted when training C-LoRA on large task sequences, as this approach suffers from scalability issues. In line with standard practices in continual learning benchmarks, we provide results for SeqLoRA, a sequential finetuning baseline without any safeguards against catastrophic forgetting. SeqLoRA experiences significant forgetting, underscoring the difficulty of maintaining performance across tasks in continual learning.
PEFT approaches with dynamic rank allocation: Parameter efficient fine-tuning techniques such as LoRA distribute trainable parameters equally across all layers, which may reduce efficiency when layers vary in their contribution towards task adaptation. To address this, adaptive low-rank adaptation strategies have emerged that dynamically adjust parameter budgets to improve fine-tuning efficiency, similar to PEARL. AdaLoRA (Zhang et al., 2023) adaptively allocates the parameter budget among weight matrices according to their importance score. The importance score is computed by parameterizing the incremental updates in the form of SV-decomposition and the singular values of unimportant updates are pruned. DyLoRA (Valipour et al., 2023) proposed a dynamic low-rank adaptation in which LoRA blocks are trained for a range of ranks instead of a single rank by sorting the representation learned by the adapter module at different ranks during training. More recent methods, such as DoRA (Mao et al., 2024) and Dynamic Rank-Selective LoRA (Lu et al., 2024), have emerged, with DoRA decomposing high-rank LoRA layers into structured single-rank components, while Dynamic Rank-Selective LoRA adaptively assigns ranks to LoRA modules based on their relevance to the current data in CL in vision language models. Although these approaches address a similar challenge as PEARL, these are not tailored for CL in custom, compact models.
B.4 Potential sensitivity of PEARL
PEARL is quite sensitive to the stability of task vector and noise contained in the resultant SVD decomposition thereof. In cases where task vectors are poorly conditioned, the dynamic rank estimation can be imprecise resulting in over or under allocation of dynamic rank. Figure 5(b) provides a depiction of resource allocation dynamics for ImageNet-R (10T) using PEARL (ViT) across three distinct runs. We note that PEARL relies solely on the proximity of a task relative to the reference task weights in parameter space. It is important to clarify that task proximity, as defined in PEARL, does not correspond to semantic similarity between tasks. Different random seeds can alter the training trajectory, thereby changing the proximity of tasks relative to the reference task. As a result, the distribution of parameter allocation can vary significantly across tasks in different runs.
B.5 Relation to Orthogonal-Subspace Methods
Orthogonal-subspace approaches (e.g., Wang et al. (2023c); Chaudhry et al. (2020)) address catastrophic forgetting in CL by enforcing geometric constraints that keep updates for new tasks orthogonal to subspaces associated with previous tasks, typically using SVD-based decompositions or orthogonalized gradients. These methods effectively reduce interference but require explicit construction, storage, and maintenance of task-specific subspaces, causing computational and memory costs to grow with the number of tasks. Moreover, they operate in the full parameter or gradient space, which can be expensive and less practical for large modern architectures. We compare against this line of work because it shares the core objective of preserving prior knowledge while providing task-specific flexibility, making it a natural conceptual baseline for evaluating PEARL. In contrast to explicit orthogonality constraints, PEARL works entirely within the parameter-efficient LoRA space and introduces dynamic rank allocation based on task similarity, allowing it to adapt capacity without building or maintaining orthogonal subspaces. This results in a more scalable and lightweight mechanism for mitigating interference, while still offering the ability to adapt to diverse tasks in a rehearsal-free continual learning setting.
Appendix C Additional Analysis
C.1 Comparison with complete parameter isolation